5L of a solution of contain 25mg of CACO3.what is its concentration in ppm???

Grade:12th Pass

1 Answers

Arun
5 years ago
Dear student
 
The concentration in ppm or the degree of hardness of water is defined as the no. of parts by mass of CaCO3 present in 1ppm by mass of hard water. According to the question 25mg of CaCO3 is present in 5L of water i.e. 5000ml of water which is equal to 5000g of water( since the density of water is 1g/cc=1g/ml, since 1cc=1ml). This implies that in 5000g of water 25mg of CaCO3 is present, which implies that in 10^6 gof water, [25mg*(10^6)]/5000 = 5ppm
 
 
Regards
Arun